OpenStudy (anonymous):

4 sin θ cos θ + 2 sin θ − 2 cos θ − 1 = 0 Find all solutions of the equation

OpenStudy (anonymous):

So far I have: \[2\sin \theta=1\] \[\sin \theta=\frac{ 1 }{ 2 }\] \[2\cos \theta=-1\] \[\cos \theta=-\frac{ 1 }{ 2 }\] Not sure what do with it next

OpenStudy (anonymous):

\[4\sin \theta \cos \theta +2\sin \theta -2\cos \theta -1 = 0\] \[2(2\sin \theta \cos \theta) +2\sin \theta -2\cos \theta -1 = 0\] \[2(\sin 2\theta) +2\sin \theta -2\cos \theta -1 = 0\]

OpenStudy (perl):

you can solve this by grouping

OpenStudy (perl):

4 sin t cos t + 2 sin t − 2 cos t − 1 = 0 2sin t ( 2 cos t + 1) - ( 2 cos t + 1 ) = 0

OpenStudy (perl):

you cant type in this program

OpenStudy (anonymous):

Haha yeah I didn't get any of that :P

OpenStudy (perl):

4 sin t cos t + 2 sin t - 2 cos t - 1 = 0 2sin t ( 2 cos t + 1) - ( 2 cos t + 1 ) = 0

OpenStudy (perl):

does that make sense so far?

OpenStudy (anonymous):

Yes it does

OpenStudy (perl):

4 sin t cos t + 2 sin t - 2 cos t - 1 = 0 2sin t ( 2 cos t + 1) - ( 2 cos t + 1 ) = 0 2sin t ( 2 cos t + 1) - 1*( 2 cos t + 1 ) = 0 (2 sint - 1 ) ( 2 cos t + 1) = 0 Using zero product property 2 sin t -1 = 0 , 2cos t + 1 = 0

OpenStudy (anonymous):

2sinθ=1 sinθ=1/2 2cosθ=−1 cosθ=−1/2 But then how do I get it so it's like 2(Pi)k+something(Pi)/something If that makes sense.

OpenStudy (anonymous):

Here I'll attach a screen shot of the example problem. Left side of picture is example problem. Green boxes have the correct answers in them.